consumer consumes only two goods for the consumer to be in equilibrium. why must rate of substitution (MRS) between the two goods must be equal to the ratio on prices of these two goods. IS it enough to ensure the equilibrium?
1
Expert's answer
2017-11-27T11:47:07-0500
Marginal rate of substitution of good or service Y for good or service X (MRSxy) is also equivalent to the marginal utility of X over the marginal utility of Y: MRSxy = MUx/MUy.
Therefore, when the consumer is choosing his utility maximized market basket on his budget line, MUx/MUy = Px/Py.
So, consumer's utility is maximized, when MRSxy = Px/Py.